# Detecting charge transfer at defects in 2D materials with electron ptychography

### v4

@@ Abstract
To conclude, we have experimentally detected charge transfer in monolayer WS2 including at its defects, via the influence of the electron distribution in the material on the phase of the probe electrons using electron ptychography and our parameter-based quantification method. Post-collection aberration correction is crucial to obtain accurate phases for charge-transfer measurements, and the convergence angle is also an important parameter, as it defines which spatial frequencies are transferred. Our results indicate that the frequencies at which charge transfer is detectable conveniently include those used for atomic-resolution imaging. Simulations based on first-principles charge redistribution due to bonding lead to an excellent agreement in the pristine case. For defective areas, we observe a notable phase increase of the metal site close to chalcogen vacancy sites, which significantly increases with the number of vacancies involved, as shown both theoretically and experimentally. Our study thus presents a significant advance in detecting atomic-scale charge transfer at defects.
@@
Supplemental Material
Supplemental Material A Methods
A.1 Sample preparation
A.2 Transmission electron microscopy
STEM experiments were conducted on a ThermoFisher Themis Z instrument equipped with a Timepix3 camera at 60 keV. The probe convergence angle was set to 20 or 30 mrad and the beam current was ca. 2 pA. Typical 4D datasets included 10241024 probe positions with a dwell time of 1–5 s. The list of events provided by the Timepix3 was converted to 4D data using an in-house code. The detector sampling was chosen to be approx. 30–40 pixels for the bright field disk and the real-space sampling was ca. 0.1–0.15 Å per pixel. The ADF detector semi-angular range was 70 to 220 mrad. To increase the signal-to-noise ratio, several frames (up to 10) where aligned and averaged. Typical resulting total doses were approx. Å.
A.3 Density functional theory
For DFT, we used the projector-augmented wave method in the open-source package GPAW [29]. The hexagonal unit cell of WS was made orthogonal, its size was optimized, and then used to create 531 supercells of WS (a total of 90 atoms) with 8.9 Å of vacuum between the periodic images of the layers (perpendicular cell size of 12 Å). To obtain relaxed structures for the multislice simulations, the atomic positions were optimized using a plane-wave basis with a cutoff energy of 500 eV and 331 k-points to sample the Brillouin zone until residual forces were below 0.02 eV/Å. To create models with defects, we removed one or more S atoms, and re-relaxed the atomic positions.
Charge transfer between the W and S columns was estimated via Bader partitioning [30] of the all-eletron charge density using the grid-based implementation of Henkelman and colleagues [31]. The resulting atomic charges for the pristine system were compared to those calculated for atoms surrounding the S vacancy.
A.4 Multislice simulations
Thermal diffuse scattering was included by running DFT-based velocity-Verlet molecular dynamics (MD) with a timestep of 2 fs, initialized with a Maxwell-Boltzmann temperature of 300 K and thermalized for 100 timesteps. Ten MD frozen-phonon snapshots were generated by saving the converged electron density every 50 timesteps, and datasets averaged over the ensemble. Although more images would likely be required for full convergence of an atomically thin material, the cost of DFT/MD for large supercells is high and we do not believe this affects our conclusions.
4D-STEM simulations were carried out using the multislice approach as implemented in the open-source package abTEM [18]. All simulation parameters were set to the experimental parameters, including the convergence angles and the dose per area simulated as Poisson noise of the diffraction patterns. To estimate statistical variation caused by the limited dose, multiple randomly noised datasets could be analyzed as indicated in the main text. The reciprocal-space sampling was 0.06 Å. For ADF, we note that DFT and IAM potentials produce essentially identical contrast.
A.5 Single-sideband ptychography
SSB ptychography was performed with the open-source PyPtychoSTEM package [34], using either the experimental or simulated 4D datasets as input. The step size was chosen to be 0.1–0.15 Å per pixel (matched in the SSB reconstruction) and the voltage was 60 kV. The convergence angle was set to 20 or 30 mrad. Post-collection aberration correction was applied using SVD to identify the residual aberrations which were then counteracted. As an example, for the case of the data shown in Fig. 1 of the main text, the SVD algorithm identified residual aberrations of C10 = -8.824 nm, C12a = 3.019 nm, C12b = 7.824 nm, C21a = -568.889 nm, C21b = -731.313 nm, C23a = 483.886 nm, C23b = 14.429 nm, C30 = 43.45 m, C32a = -2.129 m, C32b = -15.737 m, C34a = -1.167 m, C34b = -33.608 m. SFig. 1 shows the uncorrected, calculated and compensated double disk-overlap phases at different frequencies in top, middle and bottom, respectively.
A.6 Kernel method for assigning phases to atomic sites
The phases were extracted by an optimization method where atomic phases are fitted to the target image [16]. First, an initial atomic model that represents the experimental configuration is aligned with the target image. The model is converted to a point potential, which is convoluted with a kernel calculated by the contrast transfer function of SSB. The resulting simulation is compared with the target image and their correlation is maximized by optimizing the alignment between the model and the experimental image, width of the kernel and the intensities of the point potential. For the experimental data, the sample tilt is also optimized. Correlations were above 99% for the simulated data and above 93% for the experimental data.
Supplemental Material B Additional data
B.1 Effect of charge transfer on phase shifts
@@
In pristine WS2, the less electronegative tungsten atoms donate 1.36 electrons to the neighboring sulfur atoms. Using DFT-based potentials for single W atom SSB ptychography simulations, we can study the effect of charge removal or addition on the SSB phase, shown in SFigs. 2 and 3. The left panels show an isolated W atom or S2 column, which correspond to an IAM model. The middle panels show the phase for these models with an additional / reduced charge equal to 1.36 elementary charges, i.e. to a removal / addition of 1.36 electrons, corresponding to the magnitude of charge transfer in WS2. As indicated by the difference between the two plotted in the right columns, the phase change is up to 8% for the W and -5% for S2. In practise, the differences are smaller as charge is not completely removed into a uniform compensating background charge as required by DFT, but rather transferred to nearby sites. This example corresponds to perfect ionicity of the chemical bond, while in practise any covalency reduces the contrast. The crucial point, however, is that due to its effect on screening the nucleus, charge transfer can result in an order of magnitude larger change in the phase than in the potential itself, which only changes by 0.5%. This larger change in the phase substantially aids the detection of charge transfer.
B.2 IAM parametrization
@@
B.3 Phase uncertainty due to limited dose
Experimentally, sufficient precision is either achieved with high doses, combining measurements from multiple identical atomic sites, or both. Dose is the most limiting factor when it comes to phase-difference measurements, thus averaging over multiple sites is often unavoidable. The amount of electron exposure of a specimen until damage occurs depends on the material and typical doses for 2D materials range from 10–10 Å. The difference between the S2/W ratios have to be sufficiently separated so that their uncertainties are not overlapping. The error bars, which are calculated by the variation of phases due to noise, depend on the dose as well as on the number of atoms analysed. Lower statistics are desirable when the charge transfer of individual atoms are determined rather than the charge transfer in an ensemble of atoms. This is crucial when analysing defects as their configurations are often unique.
To estimate the dose-dependent precision of the phase extraction of charge transfer due to bonding for WS2, we show in SFig. 5 the effect on the quantified S2/W phase ratio for the IAM and DFT potentials using first just three pairs of sites. Since a ratio is calculated in this analysis, the relative errors between both atomic sites are summed for the total relative error. The uncertainties are sufficiently small to distinguish the IAM (no bonding) and DFT (bonding) simulations for doses of Å and above for three measurements. The error decreases with the square root of the number of sites, . Thus for instance with 10 pairs of sites, the influence of charge transfer is significant compared to the uncertainty at doses at least as low as Å.
B.4 Phase shifts at defective sites
@@
SFig. 7a shows a simulation of the defective area that is used for the analysis in Fig. 3 of the main text containing several monovacancies. SFig. 7b shows another example of a small region of the experimental SSB image of a defective region. The experimental phases from all the sites in SFig. 7b are extracted and shown in the corresponding histograms on the right. For this specific region, the S vacancies, the pristine sites (S and W) and the W close to the S vacancies (W@Svac) are identified manually. As one can see, the W@Svac sites have a higher phase than the corresponding pristine W. This is a result of the charge transfer at the defective sites, as discussed in the main text. The magnitude of the shift is related to the amount of charge transfer.
@@
SFig. 8 further shows SSB image simulations based on the DFT and IAM potentials as well as the difference between the IAM and DFT SSB images that result from charge transfer for an increasing number of defects. The first three rows show one to three monovacancies (MVs) surrounding a W site, the last three also include divacancies (DVs). Note that the DVs are based on removing additional S atoms from the MV sites, meaning there is a systematic increase of the density of defects from left to right. The contrast is kept constant, with a set range of displayed phase values for each row of images in the figure. It is clear that the difference increases, specifically at the W site, with a higher density of defects.
@@
Finally, SFig. 9 shows the IAM and DFT projected potentials and their differences for the same defect configurations as in SFig. 8. The same trend as in the SSB images is observed: with a higher defect density (left to right), the DFT potential is increasing at the W site next to the S vacancies, compared to the IAM potential. This demonstrates that the phases are directly related to the potentials and the changes in them are caused by bonding and charge transfer.
To elucidate the phase increase at the W site with increasing density of defects, we calculated the 3D pseudo valence electron density for the pristine WS2, the 3MV case, the 3DV case, and an isolated W atom. From the 3D volume, we calculated 2D electron density maps of a single slice of 0.07 Å thickness in top views and side views at the position of the W atom close to the defects. At this slice, the electron density of the S atoms are negligible in the top view. In all cases, the electron density of the 3DV configuration is lower, with the biggest difference in the side view. This is in agreement with the phase contrast we observe at different defect densities: the lower screening of the smaller electron density of W in the 3DV case leads to a higher phase.

### v5

@@ 3 Conclusions
To conclude, we have experimentally detected charge transfer in monolayer WS2 including at its defects, via the influence of the electron distribution in the material on the phase of the probe electrons using electron ptychography and a parameter-based quantification method. Post-collection aberration correction is crucial to obtain accurate phases for charge-transfer measurements, and the convergence angle is also an important parameter, as it defines which spatial frequencies are transferred. Our results indicate that the frequencies at which charge transfer is detectable conveniently include those used for atomic-resolution imaging. Simulations based on first-principles charge redistribution due to bonding lead to an excellent agreement in the pristine case. For defective areas, we observe a notable phase increase of the metal site close to chalcogen vacancy sites, which significantly increases with the number of vacancies involved, as shown both theoretically and experimentally. Our study thus presents a significant advance in detecting atomic-scale charge transfer at defects.
4 Methods
4.1 Sample preparation
4.2 Transmission electron microscopy
STEM experiments were conducted on a ThermoFisher Themis Z instrument equipped with a Timepix3 camera at 60 keV. The probe convergence angle was set to 30 mrad unless noted otherwise, and the beam current was ca. 2 pA. Typical 4D datasets included 10241024 probe positions with a dwell time of 1–5 s. The list of events provided by the Timepix3 was converted to 4D data using an in-house code. The detector sampling was chosen to be approx. 30–40 pixels for the bright field disk and the real-space sampling was ca. 0.1–0.15 Å per pixel. The ADF detector semi-angular range was 70 to 220 mrad. To increase the signal-to-noise ratio, several frames (up to 10) where aligned and averaged. Typical resulting total doses were approx. Å2.
4.3 Density functional theory
For DFT, we used the projector-augmented wave method in the open-source package GPAW [32]. The hexagonal unit cell of WS2 was made orthogonal, its size was optimized, and then used to create 531 supercells of WS2 (a total of 90 atoms) with 8.9 Å of vacuum between the periodic images of the layers (perpendicular cell size of 12 Å). To obtain relaxed structures for the multislice simulations, the atomic positions were optimized using a plane-wave basis with a cutoff energy of 500 eV and 331 k-points to sample the Brillouin zone until residual forces were below 0.02 eV/Å. To create models with defects, we removed one or more S atoms, and re-relaxed the atomic positions.
Charge transfer between the W and S2 columns was estimated via Bader partitioning [33] of the all-electron charge density using the grid-based implementation of Henkelman and colleagues [34]. The resulting atomic charges for the pristine system were compared to those calculated for atoms surrounding the S vacancy.
4.4 Multislice simulations
Thermal diffuse scattering was included by running DFT-based velocity-Verlet molecular dynamics (MD) with a timestep of 2 fs, initialized with a Maxwell-Boltzmann temperature of 300 K and thermalized for 100 timesteps. Ten MD frozen-phonon DFT-potential snapshots were generated by saving the converged electron density every 50 timesteps, and datasets averaged over the ensemble. Although more images would likely be required for full convergence of an atomically thin material, the cost of DFT/MD for large supercells is high and considering the small effect of TDS on the phase ratios, this does not affects our conclusions.
4D-STEM simulations were carried out using the multislice approach as implemented in the open-source package abTEM [21]. All simulation parameters were set to the experimental parameters, including the convergence angles and the electron dose per area simulated as Poisson noise of the diffraction patterns. To estimate statistical variation caused by the limited dose, multiple randomly noised datasets were analyzed as indicated in the main text. The reciprocal-space sampling was 0.06 Å-1. For ADF, we note that DFT and IAM potentials produce essentially identical contrast.
4.5 Single-sideband ptychography
SSB ptychography was performed with the open-source PyPtychoSTEM package [37], using either the experimental or simulated 4D datasets as input. The step size was 0.1–0.15 Å per pixel depending on the dataset and the voltage was 60 kV. The convergence angle was set to 20 or 30 mrad. Post-collection aberration correction was applied using singular value decomposition to identify the residual aberrations which were then counteracted. More details are given in the Supplementary Material.
4.6 Kernel method for assigning phases to atomic sites
The phases were extracted by an optimization method where atomic phases are fitted to the target image [17]. First, an initial atomic model that represents the experimental configuration is aligned with the target image. The model is converted to a point potential, which is convoluted with a kernel calculated by the contrast transfer function of SSB. The resulting simulation is compared with the target image and their correlation is maximized by optimizing the alignment between the model and the experimental image, the width of the kernel and the intensities of the point potential. For the experimental data, the sample tilt is also optimized. Correlations were above 99% for the simulated data and above 93% for the experimental data.
Acknowledgement
@@
Supplementary Material
S7 Post-acquisition aberration correction
Correcting the residual aberrations in phase images is crucial to determine the phase changes imposed by the sample accurately. In SSB, the double-disk overlaps in probe reciprocal space can be used to analyze any aberrations present. SFig. 1 shows examples of the double-disk overlaps from one of our datasets for four different spatial frequencies. The top row shows the experimental phases, the middle row the calculated phases with the identified aberrations and the bottom row the compensated phases, which are much more flat. The phase of the double-disk overlaps should be flat in the absence of aberrations. The residual aberrations identified by singular value decomposition (SVD) in the data shown in Fig. 1 are listed in Table 1. Note that we used a fifth order electron optical aberration corrector, which makes the magnitudes of the lower orders higher to compensate the higher-order aberrations.
C10 | 3.923 nm |
C12a | 4.494 nm |
C12b | 0.719 nm |
C21a | 425.722 nm |
C21b | -538.706 nm |
C23a | -29.749 nm |
C23b | -173.962 nm |
C30 | -18.358 m |
C32a | -23.239 m |
C32b | -30.227 m |
C34a | 2.782 m |
C34b | -0.213 m |
C41a | -0.491 mm |
C41b | 0.322 mm |
C43a | -0.292 mm |
C43b | 0.369 mm |
C45a | -0.083 mm |
C45b | -0.134 mm |
C50 | 7.208 mm |
C52a | 24.914 mm |
C52b | 34.317 mm |
C54a | -9.11 mm |
C54b | 10.464 mm |
C56a | -5.105 mm |
C56b | 8.305 mm |
S8 dCoM based charge density maps
SFig. 2 illustrates the benefits of ptychographic phase imaging over direct charge-density imaging based on the CoM data. The dCoM charge-density map is not only affected by aberrations, but is also far noisier at a given dose. A Gaussian filter reveals the atomic structure, but the strong variation over the atomic sites in the filtered images is still clearly too high to quantify the charge transfer. The ptychographic phase image with post-collection aberration correction on the other hand is far clearer and sufficiently sensitive to detect the charge transfer.
S9 Effect of charge transfer on phase shifts
@@
In pristine WS2, the less electronegative W atoms donate 1.36 electrons to the neighboring S. Using DFT potentials for W atom SSB ptychography simulations, we can study the effect of charge removal or addition on the SSB phase, as shown in SFigs. 3 and 4. The left panels show an isolated W atom or S2 column, which corresponds to an IAM model. The middle panels show the phase for these models with an additional / reduced charge equal to 1.36 elementary charges, i.e. to a removal / addition of 1.36 electrons, corresponding to the charge transfer in WS2. As indicated by the difference between the two plotted in the right columns, the phase change is up to 8% for the W and -5% for S2. In practice, the differences are smaller as charge is not completely removed into a uniform compensating background charge as required by DFT, but rather transferred to nearby sites. This example corresponds to perfect ionicity of the chemical bond, while in practice any covalency reduces the contrast. The crucial point, however, is that due to its effect on screening the nucleus, charge transfer can result in an order of magnitude larger change in the phase than in the potential itself, which only changes by 0.5%. This larger change in the phase substantially aids the detection of charge transfer.
S10 IAM parametrization
@@
S11 Phase uncertainty due to limited dose
Experimentally, sufficient precision is either achieved with high doses, combining measurements from multiple identical atomic sites, or both. Dose is the most limiting factor when it comes to phase-difference measurements, thus averaging over multiple sites is often unavoidable. The amount of electron dose a specimen can handle until damage occurs depends very much on the material and typical doses for imaging 2D materials range from 104–106 Å-2. The difference between the S2/W ratios have to be sufficiently separated so that their uncertainties are not overlapping. The error bars, which are calculated by the variation of phases due to noise, depend on the dose as well as on the number of atoms analysed. Thus obtaining sufficient statistical precision for configurations with small numbers of atoms is much more difficult.
To estimate the dose-dependent precision of the phase extraction of charge transfer due to bonding for WS2, we show in SFig. 6 the effect on the quantified S2/W phase ratio for the IAM and DFT potentials using first just three pairs of sites. Since a ratio is calculated in this analysis, the relative errors between both atomic sites are summed for the total relative error. The uncertainties are sufficiently small to distinguish the IAM (no bonding) and DFT (bonding) simulations for doses of Å-2 and above for three measurements. The standard error of the mean decreases with the square root of the number of sites, . Thus for instance with 10 pairs of sites, the influence of charge transfer is significant compared to the uncertainty at doses at least as low as Å-2.
S12 Phase shifts at defective sites
@@
SFig. 8a shows a simulation of the defective area that is used for the analysis in Fig. 3 of the main text containing several monovacancies. SFig. 8b shows another example of a small region of the experimental SSB image of a defective region. The experimental phases from all the sites in SFig. 8b are extracted and shown in the corresponding histograms on the right. For this specific region, the S vacancies, the pristine sites (S2 and W) and the W close to the S vacancies (W@Svac) are identified manually. As one can see, the W@Svac sites have a higher phase than the corresponding pristine W. This is a result of the charge transfer at the defective sites, as discussed in the main text. The magnitude of the shift is related to the amount of charge transfer.
@@
SFig. 9 shows SSB image simulations of structures with an increasing number of vacancies inserted around a W atom. Simulations based on both the DFT and IAM potentials as well as the difference between the corresponding SSB images are presented. The difference images show the result of the charge transfer, as the charge transfer is the only significant difference between the DFT and IAM images of a given model. The first three columns show one to three monovacancies (MVs) surrounding a W site. In the last three columns divacancies (DVs) are formed by removing the remaining S atoms from the MV sites, meaning there is a systematic increase of the density of vacancies from left to right. The image contrast is kept constant, with a set range of displayed phase values for each row of images in the figure. It is clear that the difference between the DFT and IAM increases with the vacancy density, especially at the W site around which the vacancies are introduced.
@@
SFig. 10 shows the IAM and DFT projected potentials and their differences for the same defect configurations as in SFig. 9. The same trend as in the SSB images is observed: as the vacancy density increases (left to right), the DFT potential increases at the W site next to the S vacancies compared to the IAM potential. This demonstrates that the phases are directly related to the potentials and the changes in them are caused by bonding and charge transfer.
To elucidate the origin of the increase of the phase at the W site that occurs with the increase in the density of vacancies around it, we calculated the 3D pseudo valence electron density for pristine WS2, V3S, V, and an isolated W atom. From the 3D volume, we calculated top and side view 2D electron density maps of a single slice of 0.07 Å thickness along the WS2 sheet at the position of the W atom adjacent to the vacancies, as shown in SFig. 11. In this thin slice centered on the W layer of atoms, the electron density of the S atoms is not visibly apparent as viewed from the top. In both top and and side view electron density maps, the electron density can be seen to shift away from W site as vacancies are introduced with the V configuration having the lowest density of valance electrons around the W@Svac and the greatest increase between it and the neighboring W atoms. Note that the completely isolated W atom has the greatest concentration of valance electron density around it. Therefore it is not simply the absence of S atoms around the a W atom that reduces the concentration of valence electrons around a W@Svac, but rather the charge transfer involved in a change in the bonding as the vacancies are introduced to the pristine structure. Compared to the isolated W atom, the electron density shifts away from the W atom when it is placed in the pristine structure. The electron density then shifts further away from the W atom as vacancies are introduced around it and the bonds essentially extend across longer distances. This is in agreement with the phase contrast we observe at different vacancy densities: the lower screening coming from the reduction of the electron density at a W site as the amount of vacancies increases also increases the phase at the W@Svac. Thus the phase of the W@Svac is greatest for the V configuration.